Can you check my answer? I am solving for x.
x+4 = 3x
___ = ____
-4 = 36
Answer: x=6

Found 3 solutions by checkley71, doukungfoo, edjones:
Answer by checkley71(8403)

IF YOUR PROBLEM IS:
(X+4)/-4=3X/36 THEN: CROSS MULTIPLYING YOU GET
-12X=36X-144
-12X-36X=-144
-48X=-144
X=-144/-48
X=3 ANSWER.

Answer by doukungfoo(195)

is this your problem?
x+4=3x
first move all terms with x to one side of the equation
this is done by subtracting x from both sides
x-x+4=3x-x
0+4=2x
4=2x
now divide both sides by 2
4/2=2x/2
2=x
now check the answer in the original equation
x+4=3x
2+4=(3)(2)
6=6
So the answer is x=2

Answer by edjones(8007)

(x+4)/4=3x/36
(x+4)/4=x/12
3(x+4)=x
3x+12=x
2x=-12
x=-6
If you CHECK ALL your problems by replacing x with your answer you will find your mistakes and learn a lot. you would be surprised how many of my mistakes I find checking.
(-6+4)/4=-6*3/36
-1/2=-1/2
Ed